# Effects of domain walls and chiral supercurrent in quantum anomalous Hall Josephson junctions

### v2

@@ Abstract
The intriguing interplay between topology and superconductivity has attracted significant attention, given its potential for realizing topological superconductivity. In this study, we investigate the transport properties of the chiral Josephson effect in the quantum anomalous Hall insulators (QAHIs)-based junction. We reveal a systematic crossover from edge-state to bulk-state dominant supercurrents, with a notable $0-\pi$ transition observed under non-zero magnetic flux through chemical potential adjustments. This transition underscores the competition between bulk and chiral edge transport. Furthermore, we identify an evolution among three distinct quantum interference patterns: from a $2\Phi_{0}$-periodic oscillation pattern, to a $\Phi_{0}$-periodic oscillation pattern, and then to an asymmetric Fraunhofer pattern ($\Phi_{0}=h/2e$ is the flux quantum, $h$ the Planck constant, and $e$ the electron charge). Subsequently, we examine the influence of domains on quantum interference patterns. Intriguingly, a distinctive Fraunhofer-like pattern, with periods twice those of the conventional one, emerges due to the coexistence of chiral edge states and domain wall states, even when the chemical potential is within the gap. These results not only advance the theoretical understanding but also pave the way for the experimental discovery of the chiral Josephson effect based on QAHI doped with magnetic impurities.
@@ I Introduction
Quantum anomalous Hall insulators (QAHIs) have emerged as distinctive topological materials in recent decades Haldane ; QAH2 ; QAH3 ; XDH . One of their prominent features is the presence of unidirectional chiral edge states, observable even in the absence of external magnetic fields. The QAHI states have been theoretically predicted in various materials QAHt1 ; QAHt2 ; QAHt3 and experimentally confirmed in films doped with Cr in $\rm{(Bi,Sb)_{2}Te_{3}}$ QAH1 ; QAH11 ; QAH12 ; QAH13 ; QAH14 ; QAH15 ; QAH16 ; QAH17 ; QAH18 ; QAH19 , V in $\rm{(Bi,Sb)_{2}Te_{3}}$ QAH4 ; QAH12 ; QAH21 ; QAH22 ; QAH23 ; QAH24 , $\rm{MnBi_{2}Te_{4}}$ QAH5 ; QAH31 ; QAH32 , and recently moiré superlattice systems QAH6 ; QAH7 ; QAH8 . Remarkably, the proximity of a QAHI to an s-wave superconductor (SC) was predicted to potentially realize a chiral topological superconductor within these heterostructures via the superconducting proximity effect TSC7 ; CMF1 . Therefore, QAHI/SC heterostructures have garnered considerable attention as promising platforms for realizing topological superconductors E1 ; E2 ; E3 ; TSC1 ; TSC2 ; TSC3 ; TSC4 ; TSC5 ; TSC6 ; TSC7 ; YNX ; Tanaka ; ZYF . More interestingly, when a QAHI is sandwiched by two QAHI/SC heterostructures, a chiral Josephson junction (JJ) can be formed. Investigating JJs based on QAHIs provides insights into the interplay between topology and superconductivity, offering the potential for novel phenomena and applications in quantum information processing and topological quantum computing Asano ; f5 ; YQ ; Ryota ; QC .
@@
Figure 1: (a) Schematic diagram of a QAHI-based JJ. The QAHI is partially covered by two superconducting electrodes. The central region and the superconducting region have sizes of $W\times L$ and $W\times L_{s}$, respectively. (b) Schematic diagram of CARs in real space, where Andreev pairs are constituted by electrons and holes propagating along opposing edges. (c) Schematic diagram of LARs in real space, depicting the formation of Andreev pairs primarily from edge carriers at a single edge. Additionally, Andreev pairs are present in the inner QAHI region, as a result of the bulk carriers. This particular regime results in transport being predominantly governed by edge carriers. (d) Schematic diagram of LARs in real space, where Andreev pairs formed by bulk carriers exhibit uniform distribution within the central region. In Figs. (b)-(d), we use purple (gray) lines to represent edge (bulk) carriers, respectively.
As illustrated in Fig. 1(a), the chiral JJ is formed by attaching two superconducting electrodes to the top of the QAHI. Typically, the supercurrent within the JJ is carried by Andreev bound states, a consequence of successive Andreev reflections (ARs) occurring at the two QAHI-SC interfaces. Each AR involves the conversion of an electron into a hole that occurs at the QAHI-SC interface AR , which is accompanied by the creation of a Cooper pair in the superconducting region. For example, an electron propagating along the top edge may experience a crossed (non-local) AR (CAR) process at the QAHI-SC interface in the absence of bulk carriers. This process would result in the electron changing into a hole at the bottom edge and the formation of an Andreev pair (see Fig. 1(b)). Thus, the critical Josephson current $I_{c}$ in response to the external magnetic flux $\Phi$ can be described theoretically by the periodicity of $2\Phi_{0}$ f1 ; f2 ; f3 ; f4 ; f5 ; f6 ; f7 , where $\Phi_{0}=h/2e$ is the flux quantum, $h$ the Planck constant, and $e$ the electron charge. The period is twice that of the conventional Aharonov-Bohm (AB) oscillations. However, there are two important questions that must be addressed in realistic JJs. First, during experimental measurements, bulk carriers may participate in the transport process in realistic JJs. Observing the evolution of the quantum interference pattern with respect to the ratio of bulk carriers to chiral edge carriers as the chemical potential $\mu$ is altered is desirable. Second, in the magnetically doped QAHI region, domain walls (DWs) are an inevitable presence. The effects of the DWs on the quantum interference pattern are worth studying.
@@
In this work, we systematically explore a JJ based on a QAHI with realistic magnetic domains, aiming to address the two aforementioned questions. The Josephson current $I_{s}$ is calculated by the recursive Green-function method. First, we look into the evolution of the current-phase relation as a function of the chemical potential $\mu$. A notable $0-\pi$ transition is observed under a non-zero magnetic flux, when the chiral edge states penetrate into the bulk region. Then, we display the critical Josephson current $I_{c}$ as a function of the magnetic flux $\Phi$ for various chemical potentials $\mu$. We find three distinct quantum interference patterns: a $2\Phi_{0}$-periodic oscillation pattern, a $\Phi_{0}$-periodic oscillation pattern, and an asymmetric Fraunhofer pattern. These three interference patterns indicate different AR processes as shown in Figs. 1(b)-(d). Further, we examine the effects of DWs on the $2\Phi_{0}$-periodic oscillation pattern and asymmetric Fraunhofer pattern. We find that the asymmetric Fraunhofer pattern is very robust against the DWs, while the $2\Phi_{0}$-periodic oscillation pattern is susceptible to modest magnetic fluctuations. Remarkably, a Fraunhofer-like pattern, with periods twice those of the conventional one, emerges when the chiral edge state and magnetic domain wall states coexist at $\mu=0$. The observed Fraunhofer-like pattern is expected to extend beyond our model and could be prevalent in materials doped with magnetic impurities, suggesting potential for further experimental verification.
@@
The paper is organized as follows. We introduce the model employed in this work in Sec. II. Sec. III presents the key findings of our calculations and the corresponding remarks. More specifically, in Sec. III.1, we study the current-phase relation as a function of chemical potential $\mu$ and observe a $0-\pi$ transition when the chiral edge states at the left and right QAHI-SC interfaces overlap in space with non-zero magnetic flux $\Phi$. Sec. III.2 discusses three distinct quantum interference patterns. The effects of DWs, which are covered in Sec. III.3, inevitably occur in the magnetically doped topological insulators. We find that the $2\Phi_{0}$-periodic oscillation effect is destroyed by modest magnetic fluctuations. A Fraunhofer-like pattern is observed even when $\mu=0$. Sec. IV is the conclusion.
II Model
@@
As depicted in Fig. 1(a), the QAHI-based junction comprises three parts: two superconducting electrodes and a central region with QAHI. The $4\times 4$ low-energy effective Hamiltonian describing magnetic topological insulator thin films in the central region is given by Yu ; WJ
@@
in the basis of $\psi_{\mathbf{k}}=\left[\psi_{\mathbf{k},t\uparrow},\psi_{\mathbf{k},t\downarrow},\psi_{\mathbf{k},b\uparrow},\psi_{\mathbf{k},b\downarrow}\right]^{\mathrm{T}}$ with the wave vector $\mathbf{k}$. $\uparrow$ ($\downarrow$) represents the spin direction, and $t$ ($b$) denotes the surface states of the top (bottom) layer. The symbols $\sigma_{x,y,z}$ and $\tau_{x,y,z}$ correspond to the Pauli matrices for spin and layer, respectively. $M_{z}$ represents the exchange field along the $z$ axis induced by the ferromagnetic ordering. The term $m(\mathbf{k})=m_{0}-m_{1}\mathbf{k}^{2}$ describes the coupling between the top and bottom surface states. The system is in the QAHI phase with the Chern number $C=\mathrm{sgn}(M_{z})$ when $\left|M_{z}\right|>\left|m_{0}\right|$, while it is a normal insulator with $C=0$ when $\left|M_{z}\right|<\left|m_{0}\right|$.
@@
$t_{0}$ $=\left(m_{0}-4\frac{m_{1}}{a^{2}}\right)\tau_{x}\sigma_{0}+M_{z}\tau_{0}\sigma_{z},$
@@
In proximity to an s-wave SC, a finite pair amplitude is induced in the QAHI. The Bogoliubov-de Gennes (BdG) Hamiltonian of QAHI/SC heterostructures in Nambu space is given by $H_{\mathrm{BdG}}=\allowbreak\sum_{\mathbf{k}}\Psi_{\mathbf{k}}^{\dagger}H_{\mathrm{BdG}}(\mathbf{k})\Psi_{\mathbf{k}}/2$ with $\Psi_{\mathbf{k}}=[(\Psi_{\mathbf{k},t\uparrow},\Psi_{\mathbf{k},t\downarrow},\Psi_{\mathbf{k},b\uparrow},\allowbreak\Psi_{\mathbf{k},b\downarrow}),(\Psi_{\mathbf{-k},t\uparrow}^{\dagger},\Psi_{\mathbf{-k},t\downarrow}^{\dagger},\Psi_{\mathbf{-k},b\uparrow}^{\dagger},\Psi_{\mathbf{-k},b\downarrow}^{\dagger})]^{\mathrm{T}}$ and
$H_{\mathrm{BdG}}(\mathbf{k})$ $=\left(\begin{array}[]{cc}H(\mathbf{k})-\mu&\Delta\\ \Delta^{\dagger}&-H^{*}(-\mathbf{k})+\mu\end{array}\right),$ (3)
@@
Here $\mu$ is the chemical potential, $\varphi_{L}$ and $\varphi_{R}$ are the superconducting phases in the left and right sides of the junction. The physical properties only depend on the superconducting phase difference $\varphi=\varphi_{L}-\varphi_{R}$. $\Delta_{t}$ and $\Delta_{b}$ denote the induced superconducting pairing potentials on the top and bottom surface states, respectively. For simplicity, we set $\Delta_{t}=0.5$ and $\Delta_{b}=0$ in the following calculations, and both the QAHI and superconducting regions have dimensions $W\times L(L_{s})=100\times 40$.

### v3

# Anomalous Fraunhofer-like patterns in quantum anomalous Hall Josephson junctions

@@ Abstract
The intriguing interplay between topology and superconductivity has attracted significant attention, given its potential for realizing topological superconductivity. In the quantum anomalous Hall insulators (QAHIs)-based junction, the supercurrents are carried by the chiral edge states, characterized by a $2\Phi_{0}$ magnetic flux periodicity ($\Phi_{0}=h/2e$ is the flux quantum, $h$ the Planck constant, and $e$ the electron charge). However, experimental observations indicate the presence of bulk carriers in QAHI samples due to magnetic dopants. In this study, we reveal a systematic transition from edge-state to bulk-state dominant supercurrents as the chemical potential varies from the bulk gap to the conduction band. This results in an evolution from a $2\Phi_{0}$-periodic oscillation pattern to an asymmetric Fraunhofer pattern. Furthermore, a novel Fraunhoher-like pattern emerges due to the coexistence of chiral edge states and bulk states caused by magnetic domains, even when the chemical potential resides within the gap. These findings not only advance the theoretical understanding but also pave the way for the experimental discovery of the chiral Josephson effect based on QAHI doped with magnetic impurities.
@@ I Introduction
In recent decades, research on quantum anomalous Hall insulators (QAHIs) has flourished, driven not only by the intriguing physical phenomena they exhibit but also by their potential for various technological applications Haldane ; QAH2 ; QAH3 ; XDH . These applications include serving as platforms for realizing topological quantum computation, resistance standards, and dissipation-free interconnects QAH3 . In a QAHI, the bulk behaves as an insulator, while electrons can propagate along the edges of the sample. These edge states, termed chiral edge states, propagate unidirectionally along the edge. Consequently, QAHI is distinguished by a quantized Hall effect, even in the absence of external magnetic fields. The existence of QAHI states has been theoretically predicted in a range of materials QAHt1 ; QAHt2 ; QAHt3 and experimentally verified in Cr-doped topological insulator films for the first time QAH1 . Subsequently, QAHI has been experimentally demonstrated in various materials, including Cr-doped $\rm{(Bi,Sb)_{2}Te_{3}}$ films QAH11 ; QAH12 ; QAH13 ; QAH14 ; QAH15 ; QAH16 ; QAH17 ; QAH18 ; QAH19 , V-doped $\rm{(Bi,Sb)_{2}Te_{3}}$ films QAH4 ; QAH12 ; QAH21 ; QAH22 ; QAH23 ; QAH24 , $\rm{MnBi_{2}Te_{4}}$ QAH5 ; QAH31 ; QAH32 , and recently moiré superlattice systems QAH6 ; QAH7 ; QAH8 . Remarkably, the proximity of a QAHI to an s-wave superconductor (SC) was predicted to potentially realize a chiral topological superconductor within these heterostructures via the superconducting proximity effect TSC7 ; CMF1 . Therefore, QAHI/SC heterostructures have garnered considerable attention as promising platforms for realizing topological superconductors E1 ; E2 ; E3 ; TSC1 ; TSC2 ; TSC3 ; TSC4 ; TSC5 ; TSC6 ; TSC7 ; YNX ; Tanaka ; ZYF . More interestingly, when a QAHI is sandwiched by two QAHI/SC heterostructures, supercurrent is carried by the chiral edge states, termed a chiral Josephson junction (JJ) in our paper. Investigating JJs based on QAHIs provides insights into the interplay between topology and superconductivity, offering the potential for novel phenomena and applications in quantum information processing and topological quantum computing Asano ; YQ ; Ryota ; QC .
@@
Figure 1: (a) Schematic diagram of a QAHI-based JJ. The QAHI is partially covered by two superconducting electrodes, both composed of QAHI/SC heterostructures. The central region and the superconducting region have sizes of $W\times L$ and $W\times L_{s}$, respectively. (b) The QAHI energy spectrum with a length of $L=40a$ is calculated by using a lattice model described by Eq. 2. $\mu_{1}=0$ and $\mu_{2}=0.42\Delta$ are the chemical potentials, indicated by the green dashed lines. (c) Schematic diagram of CARs in real space, where Andreev pairs are constituted by electrons and holes propagating along opposing edges. (d) Schematic diagram of LARs in real space, where Andreev pairs formed by bulk carriers exhibit uniform distribution within the central region. In Figs. (b)-(d), we use purple (gray) lines to represent edge (bulk) carriers, respectively.
Generally, the supercurrent in JJs is generated by the Andreev reflection process, which involves the conversion of an electron into a hole at the superconductor interface AR . This process is accompanied by the creation of a Cooper pair in the superconducting region. As shown in Fig. 1(d), local Andreev reflections (LARs) occur uniformly in a conventional JJ. Thus, the critical Josephson current $I_{c}$, which is carried by bulk carriers, oscillates with the external magnetic flux $\Phi$ due to the interference of the supercurrents flowing through the junction in the presence of the magnetic flux. This oscillatory behavior gives rise to a quantum interference pattern known as the Fraunhofer pattern. The Fraunhofer pattern is characterized by a central lobe with a width of $2\Phi_{0}$ and decaying side lobes. Here, $\Phi_{0}=h/2e$ is the flux quantum, $h$ the Planck constant, and $e$ the electron charge. When time-reversal symmetry is preserved, the Fraunhofer pattern is symmetric, meaning $I_{c}(\Phi)=I_{c}(-\Phi)$.
On the other hand, in a chiral JJ depicted in Fig. 1(a), the presence of gapped bulk states inhibits the bulk supercurrent. However, supercurrents can propagate through the chiral edge states even when the chemical potential resides within the bulk gap. Due to the chiral nature of edge states, electrons on the top edge are transformed into holes propagating along the bottom edge (see Fig. 1(c)). This conversion of electrons and holes takes place in the spatially separated edges, known as crossed Andreev reflections (CAR) CAR1 ; CAR2 ; f2 ; f6 ; f8 . This process is reminiscent of the Aharonov-Bohm (AB) oscillations AB1 , where electrons traveling along different paths enclose a magnetic flux $\Phi$, causing resistance to oscillate with the magnetic flux exhibiting a periodicity of $2\Phi_{0}$ AB2 . Consequently, the critical Josephson current $I_{c}$ in response to the external magnetic flux $\Phi$ can be described theoretically by the periodicity of $2\Phi_{0}$ f1 ; f2 ; f3 ; f4 ; f5 ; f6 ; f7 ; f8 . On the other hand, experimental observations have confirmed the importance of the bulk states, induced by magnetic dopants, on the transport properties of QAHIs QAH2 ; QAH3 . However, previous studies did not consider the effects of these bulk states on the quantum interference patterns in chiral JJs. Therefore, research addressing this question is highly desirable.
@@
In this work, we systematically explore a JJ based on a QAHI with magnetic dopants, aiming to address the above question. The Josephson current $I_{s}$ is calculated by the recursive Green-function method. First, we present the critical Josephson current $I_{c}$ as a function of the magnetic flux $\Phi$ for various chemical potentials $\mu$. We find an evolution of quantum interference patterns from a $2\Phi_{0}$-periodic oscillation pattern to an asymmetric Fraunhofer pattern. The two typical interference patterns result from different Andreev reflection processes as shown in Figs. 1(c)-(d). Second, we simulate the effects of magnetic domains on the two interference patterns. We find that the asymmetric Fraunhofer pattern is very robust against the domains, while the $2\Phi_{0}$-periodic oscillation pattern is susceptible to modest magnetic fluctuations. Remarkably, a novel Fraunhofer-like pattern, with periods twice those of the conventional one, emerges when the chiral edge state and magnetic domain states coexist at $\mu=0$. The observed Fraunhofer-like pattern is expected to extend beyond our model and could be prevalent in materials doped with magnetic impurities, suggesting potential for further experimental verification.
@@
The paper is organized as follows. We introduce the model and methods employed in this work in Sec. II. Sec. III presents the key findings of our calculations and the corresponding remarks. More specifically, Sec. III.1 discusses two distinct quantum interference patterns. The effects of magnetic domains, which are covered in Sec. III.2, inevitably occur in the magnetically doped topological insulators. We find that the $2\Phi_{0}$-periodic oscillation effect is destroyed by modest magnetic fluctuations. A novel Fraunhofer-like pattern is observed even when $\mu=0$. Sec. IV is the conclusion.
II Model and methods
@@
As depicted in Fig. 1(a), the QAHI-based junction comprises three parts: two superconducting electrodes and a central region with QAHI. The $4\times 4$ low-energy effective Hamiltonian describing magnetic topological insulator thin films in the central region is given by QAHt3 ; WJ
@@
in the basis of $\psi_{\mathbf{k}}=\left[\psi_{\mathbf{k},t\uparrow},\psi_{\mathbf{k},t\downarrow},\psi_{\mathbf{k},b\uparrow},\psi_{\mathbf{k},b\downarrow}\right]^{\mathrm{T}}$ with the wave vector $\mathbf{k}$. $\uparrow$ ($\downarrow$) represents the spin direction, and $t$ ($b$) denotes the top (bottom) layer. The symbols $\sigma_{x,y,z}$ and $\tau_{x,y,z}$ correspond to the Pauli matrices for spin and layer, respectively. $M_{z}$ represents the exchange field along the $z$ axis induced by the ferromagnetic ordering. The term $m(\mathbf{k})=m_{0}-m_{1}\mathbf{k}^{2}$ describes the coupling between the top and bottom layers. The system is in the QAHI phase with the Chern number $C=\mathrm{sgn}(M_{z})$ when $\left|M_{z}\right|>\left|m_{0}\right|$, while it is a normal insulator with $C=0$ when $\left|M_{z}\right|<\left|m_{0}\right|$.
@@
$t_{0}$ $=\left(m_{0}-4\frac{m_{1}}{a^{2}}\right)\tau_{x}\sigma_{0}+M_{z}\tau_{0}\sigma_{z},$ (3)
@@
In proximity to an s-wave SC, a finite pair amplitude is induced in the QAHI. The Bogoliubov-de Gennes (BdG) Hamiltonian of QAHI/SC heterostructures in Nambu space is given by $H_{\mathrm{SC}}=\allowbreak\sum_{\mathbf{k}}\Psi_{\mathbf{k}}^{\dagger}H_{\mathrm{SC}}(\mathbf{k})\Psi_{\mathbf{k}}/2$ with $\Psi_{\mathbf{k}}=[(\Psi_{\mathbf{k},t\uparrow},\Psi_{\mathbf{k},t\downarrow},\Psi_{\mathbf{k},b\uparrow},\allowbreak\Psi_{\mathbf{k},b\downarrow}),(\Psi_{\mathbf{-k},t\uparrow}^{\dagger},\Psi_{\mathbf{-k},t\downarrow}^{\dagger},\Psi_{\mathbf{-k},b\uparrow}^{\dagger},\Psi_{\mathbf{-k},b\downarrow}^{\dagger})]^{\mathrm{T}}$ and
$H_{\mathrm{SC}}(\mathbf{k})$ $=\left(\begin{array}[]{cc}H(\mathbf{k})-\mu_{s}&\Delta\\ \Delta^{\dagger}&-H^{*}(-\mathbf{k})+\mu_{s}\end{array}\right),$ (4)
@@
Here, $\mu_{s}$ is the chemical potential of the superconduting regions, $\varphi_{L}$ and $\varphi_{R}$ are the superconducting phases in the left and right sides of the junction. The physical properties only depend on the superconducting phase difference $\varphi=\varphi_{L}-\varphi_{R}$. $\Delta_{t}$ and $\Delta_{b}$ denote the induced superconducting pairing potentials on the top and bottom layers, respectively.
The QAHI regions in Nambu space can be represented as
$H_{\mathrm{BdG}}(\mathbf{k})=\left(\begin{array}[]{cc}H(\mathbf{k})-\mu&0\\ 0&-H^{*}(-\mathbf{k})+\mu\end{array}\right)$ (5)
Here, $\mu$ is the chemical potential of the QAHI region. For simplicity, the parameters are set as follows: lattice constant $a=1$, Fermi velocity $\upsilon_{F}=1$, $\Delta_{t}=0.5$ and $\Delta_{b}=0$, $m_{1}=2\Delta_{t}a^{2}$, and $m_{0}=-0.2\Delta_{t}$ in this work and both the QAHI and superconducting regions have dimensions $W\times L(L_{s})=100a\times 40a$. In the following calculations, we use $\Delta$ instead of $\Delta_{t}$.
The Josephson current $I$ in the QAHI region can be calculated by the recursive Green-function method and expressed as M1 ; M2 ; M3
$I=-\frac{ieK_{B}T}{\hbar}\sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n}}\operatorname{Tr}\left[\hat{T}_{x}\hat{G}_{x+1,x}(i\omega_{n})-\hat{T}^{\dagger}_{x}\hat{G}_{x,x+1}(i\omega_{n})\right]$ (6)
where $e$ is the electron charge, $k_{B}$ is the Boltzmann constant, $\hbar$ is the reduced Planck constant, and $T$ is the temperature. The hopping term $\hat{T}_{x}$ in Nambu space is given by $\hat{T}_{x}=\left(\begin{array}[]{cc}t_{x}&\\ &-t_{x}^{*}\end{array}\right)$, where $t_{x}$ originates from Eq. 3. The Green’s function is expressed as $\hat{G}(i\omega_{n})=\left(i\omega_{n}-H_{BdG}-\Sigma_{L}-\Sigma_{R}\right)^{-1}$ with the Matsubara frequency $\omega_{n}=(2n+1)\pi k_{B}T$. The index $n$ in the sum represents a summation over Matsubara frequencies, where $n$ spans all integer values. Here, $\Sigma_{L(R)}$ denotes the self-energy induced by the coupling between the left (right) superconducting leads and the QAHI region, which can be calculated numerically lead .
Figure 2: Quantum interference patterns in different regimes. The critical Josephson current $I_{c}$ versus the magnetic flux $\Phi$ in the unit of $\Phi_{0}$. (a) CARs happen when only chiral edge states are included with $\mu=0$. $I_{c}$ exhibits the AB oscillation characterized by a periodicity of $2\Phi_{0}$. (b)-(e) The behavior of $I_{c}$ deviates from the $2\Phi_{0}$-periodic oscillation as $\mu$ takes on values of $0.2\Delta$, $0.42\Delta$, $0.8\Delta$, and $1.4\Delta$, respectively. (c) LARs take place in the JJs, and the $2\Phi_{0}$-periodic oscillation is destroyed. (d) The emergence of central and side lobes is a consequence of the interference pattern transitioning towards a regime dominated by bulk carriers. (f) $I_{c}$ displays the typical asymmetric Fraunhofer pattern with $\mu=2\Delta$. In this regime, the bulk carriers uniformly flow through the whole QAHI region. In all cases, the blue (orange) solid lines represent $I_{c}$ curves corresponding to $M_{z}=\pm 0.6\Delta$, $\mu_{s}=\mu$, and the temperature $T=\Delta/200$
Figure 3: The spatial distribution of current in the QAHI region [(a) and (c)] and at the QAHI-SC interface [(b) and (d)] when the total current has its maximum value is examined under varying chemical potentials: $\mu=$0 [(a),(b)], and $2\Delta$ [(c),(d)]. In all cases, $M_{z}=0.6\Delta$, $\Phi=0$, $\mu_{s}=\mu$, and the temperature $T=\Delta/200$.